40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10219     Entry time: Wed Jul 16 19:38:37 2014
Author: manasa 
Type: Summary 
Category: PSL 
Subject: AOM alignment issues and removed from beam path 

AOM removed from the beampath and PMC relocked. 

AOM alignment:

1. Measured the initial power after PMC as 1.30W and reduced it down to 130mW.
2. Checked the power in the AOM zero order transmission before touching it. For 0-1V modulation input, the power dropped from 125uW to 98.3uW.
3. Steered the mirror right before the AOM to increase AOM zero order transmission and then carefully moved the AOM around to obtain maximum power attenuation. I repeated this a few times and the maximum attenuation that I could obtain was 125uW to 89.2uW (~30% attenuation).
Although this is not the right way to align the AOM, we do not have much options with the current setup as there is not enough separation between the zero order and first order beams and the AOM is on a fixed rigid mount.
4. I tried to dump the first order beam from the AOM and it wasn't satisfactory as well. There is barely any separation between the zero order and first order beams.

PMC relocking:

1. SInce the alignment to the PMC was disturbed by moving the AOM and the steering mirror in front of it, the PMC alignment was lost.
2. I could not relock the PMC at low power or high power. Rana had to come to rescue and fixed the alignment so that we could see flashes of PMC on the trans camera (This was done by aligning refl beam to the PMC REFL PD while giving a triangular ramp to the PMC PZT voltage).
Also I should not have tried to lock the PMC at high power as I could have been steering the beam at high power to the edges of the PMC mirrors that way and burning stuff easily.
3. Before fine tuning the alignment, I decided to remove the AOM from the beam path as there needs some work done on it to make it useful.
4. I removed the AOM from the beam path and relocked the PMC. 
5. PMC is relocked with 0.79 counts in TRANS and I measured the power after PMC 1.30W

Attachment: picture showing AOM removed from the beampath.

Attachment 1: AOMremoved.jpg  114 kB  Uploaded Wed Jul 16 20:42:53 2014  | Hide | Hide all
AOMremoved.jpg
ELOG V3.1.3-